​(Break-even analysis)  Niece Equipment Rentals of Del​ Valle, Texas, has recently been approached about the prospect of purchasing a large construction crane. The crane rents for

​$530530

an hour but​ operator, fuel, insurance and miscellaneous expenses run

​$204204

an hour when the crane is in use. The company owner estimates that it will cost

​$1 comma 0401,040

a month to store and maintain the crane and the annual depreciation expense is

​$45 comma 00045,000.

a.  Calculate the accounting​ break-even number of annual rental hours needed to produce zero operating earnings from the crane​ (before taxes).

b.  Calculate the cash​ break-even point. If we ignore​ non-cash expenses such as depreciation in the​ break-even calculation, how many hours must the crane be rented in order to break even on a cash​ basis?

c.  Why do we have two different​ break-even points? What does each one tell​ you?

a.  The accounting​ break-even units of production is _______ hours. 
Answer #1

Annual Depreciation expense of the crane= $45000

Store and maintenance cost of the crane= monthly cost*number of month in year=401040*12=$4812480

Operating profit per hour when crane is being rented= (530530-204204)=$326326

Say, number of rental our needed is x

Hence, x* operating profit per hour=storage and maintenance cost+Annual depreciation

or, x*326326=(4812480+45000)

or, x=14.88 or 15 hour (Approx)

a. Hence, The accounting break even units of production is 15 hours approx.

b. If we exclude non cash expense, then

x*326326=4812480 (operating profit need to cover only store and maintenance cost)

or, x= 14.74 or 15 hour (approx)

c. As we do not have to cover depreciation expense for cash break even point, we have lesser break even point as hour.